# From LLMs to MLLMs to Agents: A Survey of Emerging Paradigms in Jailbreak Attacks and Defenses within LLM Ecosystem

### v2

@@ IV-A Questions-Answers (Q&A)
First, for question-answering datasets, Gressel et al. [37] construct a benchmark dataset containing 210 prompts. The dataset sources include academic literature, Twitter, Medium, and other online platforms, with a strong emphasis on prompt diversity. It is categorized by “strategies,” each comprising multiple “techniques,” and each technique includes five variants to account for the randomness of LLM text generation. The dataset is divided into two main types: implicit challenges (8 strategies, 33 techniques, 165 prompts), where the LLM autonomously determines whether it deviates from its assigned role, and explicit challenges (2 strategies, 9 techniques, 45 prompts), which can be automatically verified by programs. Röttger et al. [98] introduce a new test suite named XSTEST, which systematically identifies overly cautious behavior in safety alignment. XSTEST consists of 250 safe prompts and 200 unsafe prompts. The safe prompts span 10 distinct types and are carefully designed such that a well-calibrated model should accept and execute them. In contrast, the unsafe prompts involve content that should be universally rejected, enabling a comprehensive assessment of the model’s ability to balance safety and usability.
The StrongREJECT [100] dataset contains carefully designed prompts that compel the target model to provide specific, harmful content. They also develop an automatic evaluator that quantifies the harmfulness of the model’s responses, thus measuring jailbreak effectiveness. Notably, the StrongREJECT evaluator achieves state-of-the-art agreement with human judgments in assessing jailbreak success. Furthermore, they find that existing evaluation methods significantly overestimate jailbreak success compared to StrongREJECT and human evaluations. To explain this discrepancy, Souly et al. describe a novel and unexpected phenomenon: when a jailbreak successfully bypasses the model’s safety fine-tuning, it often degrades the model’s overall performance, revealing potential impacts of jailbreaks on capability. Zhang et al. [95] propose SafetyBench, a comprehensive benchmark for evaluating LLM safety. SafetyBench includes 11,435 multiple-choice questions covering seven major safety-related categories. The majority of these questions are generated by transforming existing dataset samples, complemented by manually crafted safety problems to ensure diversity and rigor. Importantly, SafetyBench includes both Chinese and English data, allowing for bilingual safety evaluations. Jin et al. [96] construct a real-world dataset specifically designed for jailbreak prompts. It includes 666 jailbreak prompts and 390 harmful questions across 13 high-risk scenarios, such as illegal activities, hate speech, and malware generation. The data is sourced from Reddit, Discord, various websites, and open-source datasets, ensuring diversity and realism. To improve quality, they curate a ground truth dataset by selecting the most effective responses for each question and use BERT embeddings to compute answer similarity for evaluating response effectiveness.
In addition, Gong et al. [71] construct SafeBench, a safety evaluation benchmark that includes 500 harmful questions across 10 key AI safety constraint topics, such as illegal activity, hate speech, and malware generation. The dataset is built in two phases: first, sensitive topics are identified based on OpenAI and Meta usage policies; second, GPT-4 is used to generate questions, which are then manually filtered to ensure they violate AI safety policies. For scalability, Gong et al. [71] also create SafeBench-Tiny, a smaller subset with 50 randomly selected questions. Banerjee et al. [102] develop the TECHHAZARDQA dataset, which includes 7,745 harmful questions covering seven technical domains, including biotechnology, cybersecurity, and finance. The dataset construction process involves using an unsafely fine-tuned Mistral-V2 model to generate a large number of potentially harmful questions, filtering those answerable by text or pseudocode. They then manually review the questions to ensure they elicit unsafe model behavior while discarding harmless or irrelevant items. The dataset is used to evaluate LLM safety and vulnerability under zero-shot, zero-shot CoT, and few-shot prompting strategies. Wang et al. [101] introduce a three-level risk taxonomy ranging from mild to extreme risk. Based on this framework, they collect at least 10 prompts for each category and construct a risk detection dataset comprising 939 prompts, all of which should not be executed by a safe model. The fine-grained taxonomy helps reveal specific vulnerabilities that LLMs must prioritize mitigating. Liu et al. [97] construct a training dataset containing 1.2 million publicly available image-text pairs, covering various academic tasks such as visual question answering (VQA). The dataset construction process involves extracting visual features using CLIP-ViT-L-336px and optimizing them via an MLP projection. They also incorporate VQA data specific to academic tasks and introduce prompt formatting to enhance model performance across tasks.
@@ IV-B Instructions
For instruction-style datasets, Qiu et al. [99] propose a benchmark to evaluate the safety and robustness of LLMs, emphasizing the importance of achieving a balance between the two. They introduce a latent jailbreak prompt dataset containing malicious instructions embedded within seemingly benign tasks, such as translation (where the text to be translated contains the malicious instruction). To facilitate in-depth analysis of safety and robustness, the researchers design a hierarchical annotation framework and systematically analyze LLM performance across several dimensions, including explicit benign instructions, word substitutions (e.g., verbs in benign prompts, target groups in malicious prompts, cue words in benign prompts), and instruction placement. SAP (Semi-Automatic Attack Prompts) [89] is an attack prompt dataset specifically designed for LLM safety evaluation and defense research. It consists of multiple versions: SAP5, SAP10, SAP20, SAP30, and SAP200, with SAP200 comprising 1,600 carefully crafted attack prompts covering sensitive topics such as fraud, politics, pornography, race, religion, suicide, terrorism, and violence, ensuring wide applicability in multi-aspect safety evaluations. SAP construction combines manual and automated approaches. Initially, high-quality, manually created prompts are collected from prior research and public resources, including online jailbreak prompt libraries such as those for ChatGPT. Researchers then use GPT-3.5-Turbo-0301 as the attack model to automatically expand the dataset via in-context learning. Specifically, the model is prompted with several high-quality examples and asked to imitate them to generate new attack prompts. To improve quality, a Chain-of-Thought-like strategy is adopted, requiring GPT-3.5-Turbo-0301 to explain the harmful nature of each example prompt, thereby guiding the generation of prompts with attack characteristics. After generation and evaluation, high-quality prompts are filtered into the dataset, which is iteratively expanded and optimized via repeated in-context learning, enhancing its effectiveness in LLM safety assessment.
Subsequently, to facilitate research on LLM agent misuse, Andriushchenko et al. [42] introduce a new benchmark framework called AgentHarm. This benchmark includes 110 clearly defined harmful agent tasks (based on 110 core behaviors), forming a dataset of 440 tasks. The tasks span 11 harm categories, including fraud, cybercrime, self-harm, harassment, sexual content, copyright infringement, drug-related content, information leakage, hate speech, violence, and terrorism. In addition to assessing whether the model refuses harmful agent requests, high scores on AgentHarm also require jailbreak agents to maintain their functional capabilities post-attack to complete multi-step tasks. JBB-Behaviors [92] is a component of the JailbreakBench benchmark, specifically designed to evaluate the safety and defense capability of LLMs under jailbreak attacks. The dataset includes 100 harmful behaviors and their corresponding 100 benign counterparts, covering categories such as harassment, malware, physical harm, financial harm, fraud, misinformation, adult content, privacy violations, misuse of expert advice, and government interference. The data sources include original contributions by the authors (55%), TDC/HarmBench (27%), and AdvBench (18%). During construction, the researchers apply jailbreak methods such as PAIR, GCG, and JBC across multiple LLMs (e.g., Vicuna, Mistral, LLaMA) to generate jailbreak prompts and create matched benign behaviors for each harmful behavior, ensuring fair testing. This dataset provides a standardized evaluation framework for LLM safety research, enabling more systematic and reproducible comparisons across different attack and defense strategies.
@@ IV-C Harmful Sentences / Images
For datasets in the form of harmful sentences or images, Zou et al. [81] design a novel benchmark, AdvBench, based on two different settings. AdvBench consists of 500 strings that reflect a wide range of harmful or toxic behaviors, including profanity, depictions of violence, threats, misinformation, discrimination, cybercrime, and dangerous or illegal advice. The attacker’s objective is to craft specific inputs that induce the model to accurately generate these strings. The string lengths range from 3 to 44 tokens, with an average of approximately 16 tokens after tokenization using the LLaMA tokenizer. AdvBench also includes 500 instructions corresponding to harmful behaviors, with topics aligned with those in the harmful string setting. The attacker aims to find a universal adversarial string that triggers the model to perform as many harmful behaviors as possible when executing these instructions. Through iterative updates, AdvBench has now expanded to 574 harmful strings and 520 harmful instructions.
Subsequently, Niu et al. [61] categorize the harmful behaviors in AdvBench into eight semantic categories (e.g., “bombs or explosives,” “drugs,” “self-harm and suicide,” etc.) and retrieve 30 semantically relevant images for each category from the internet. They then pair each harmful behavior with a corresponding image to construct the multi-modal dataset AdvBench-M, which is used to evaluate the jailbreak capabilities of MLLMs. To facilitate the safe deployment of Chinese LLMs, Sun et al. [93] develop a safety evaluation benchmark for Chinese LLMs, named SafetyPrompts. This benchmark assesses the overall safety performance of LLMs along two dimensions: eight typical safety scenarios and six more challenging instruction-based attacks. The evaluation follows a straightforward process, wherein test prompts are provided and the safety of the model’s responses is assessed. In terms of methodology, the benchmark leverages the strong capabilities of LLMs and optimizes them, through reward-based prompting, into effective safety evaluators. Liu et al. [94] automatically compile a dataset of 5,040 text-image pairs using stable diffusion and layout techniques, covering 13 harmful scenarios. Li et al. [33] introduce RTVLM, a dataset that contains 750 harmful instructions across five representative categories of harmful content: violence, financial crimes, privacy violations, self-harm, and animal abuse. The dataset is constructed as follows: first, the researchers use GPT-4 to generate 50 keywords for each harmful scenario. Based on these keywords, they then generate three distinct harmful instructions. To enhance the attack’s effectiveness, they search Google using the keywords to retrieve five related images and apply CLIP ViT-L/14 to filter the images most semantically aligned with the keywords, ensuring strong relevance between visual and textual content.
@@ V Evaluation Metrics
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=root-line-color, rounded corners, align=center, text centered, minimum width=5em, edge path= [draw=edge-color, line width=1pt] (!parent.east) – +(5pt,0) —- (.child anchor); , s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=1pt, par/.style=rotate=0, child anchor=north, parent anchor=south, anchor=center, minimum height=3em, , where level=1 draw=child-one-line-color, text width=20em, minimum height=3em, font=, , where level=2 draw=child-five-line-color, minimum height=3em, font=, , [ Evaluation Metrics, par, for tree=fill=root-color, [ Human Evaluation §$\S$ V-A, for tree=fill=child-one-color, [ xx[36, 31, 35, 30, 90, 74, 76, 80]xx, for tree=fill=child-five-color, ] ] [ Perspective API Evaluation §$\S$ V-B, for tree=fill=child-two-color, [ xx[22, 28, 68, 90, 31]xx, for tree=fill=child-five-color, ] ] [ LLM Evaluation §$\S$ V-C, for tree=fill=child-three-color, [ xx[26, 25, 22, 31, 34, 33, 35, 28, 90, 29, 82, 89, 37, 74, 91, 40, 80, 84]xx, for tree=fill=child-five-color ] ] [ Keyword Dictionary Evaluation §$\S$ V-D, for tree=fill=child-four-color, [ xx[26, 90, 29, 82]xx, for tree=fill=child-five-color ] ] [ Custom Evaluation §$\S$ V-E, for tree=fill=child-six-color, [ xx[41, 43, 27, 85, 75, 81, 30, 79, 38, 68, 77, 86, 78, 39]xx, for tree=fill=child-five-color ] ] ] ]
@@ V-A Human Evaluation
This method involves systematic review of model responses by experts or professional annotators, focusing on whether the content exhibits jailbreak behaviors. The assessment includes grading the severity, executability, and potential societal impact of such behaviors. Researchers such as Yu et al. [30] and Shayegani et al. [31] predefine multiple scoring levels and provided systematic training for evaluators. During the evaluation process, each harmful text is independently scored by four experts. If all experts unanimously classify the text as harmful, it is directly categorized as a successful jailbreak case. If three experts deem it harmful, an additional senior reviewer is invited for a final decision to maximize evaluation accuracy and fairness.
Limitations: While this approach offers significant advantages in evaluation accuracy, it suffers from high costs, poor scalability, and other practical constraints, making it difficult to adapt to large-scale evaluation needs.
@@ V-B Perspective API Evaluation
This method leverages the quantitative analysis capabilities of Perspective API to automatically score text based on toxicity, insultiveness, hatefulness, and other harmful attributes [31, 28]. Specifically, a score above 0.7 indicates the presence of clearly harmful content; scores between 0.3 and 0.7 suggest a high likelihood of harmful information; and scores below 0.3 are generally considered benign.
Limitations: This approach has shortcomings in cross-language adaptability and contextual understanding, particularly struggling with nuanced language phenomena such as metaphors and sarcasm, which may lead to misclassification. Additionally, for longer texts, Perspective API tends to assign lower harmfulness scores, compromising evaluation accuracy.
@@ V-C LLM Evaluation
This approach typically employs high-performance LLMs (e.g., GPT-4, Claude) as the core evaluators, leveraging techniques such as prompt engineering to systematically analyze the harmfulness, policy-violating severity, and jailbreak effectiveness of generated text. For instance, Some researchers [25, 26, 90] instruct GPT-4 to rate the harmfulness of response texts on a scale of 1–10, where 10 indicates explicitly harmful content and is considered a successful jailbreak.
Limitations: Although this method demonstrates strong robustness in automated evaluation, in practice, the LLM may refuse to generate ratings for text containing harmful content, compromising the stability and feasibility of the assessment.
@@ V-D Keyword Dictionary Evaluation
This method employs pattern-matching techniques to rapidly detect model outputs based on predefined sensitive word libraries and keyword lists, making it the fastest evaluation approach currently available. Common keyword dictionaries typically include terms such as “I’m sorry”, “Sorry”, “I apologize”, “I’m just”, “can’t”, “I do not”, and “must” [26, 90]. If any of these dictionary terms appear in the model’s response, the jailbreak attempt is considered unsuccessful.
Limitation: The key challenge lies in constructing a well-designed dictionary that comprehensively covers the model’s refusal patterns to ensure evaluation accuracy and thoroughness.
@@ V-E Custom Evaluation
To address specific research needs, researchers can design tailored evaluation schemes by integrating rule-based matching, contextual analysis, and interactive feedback to construct targeted assessment frameworks. Specifically, we categorize these custom evaluation methods into three groups, depending on whether the jailbreak attacks target LLMs, MLLMs, or Agents.
V-E1 Jailbreak Evaluation for LLMs
@@
Yu et al. [30] propose two metrics for evaluating jailbreak effectiveness: Expected Maximum Harmfulness (EMH) and Jailbreak Success Rate (JSR). EMH aims to measure the maximum potential harm caused by jailbreak inputs under the worst-case scenario. Specifically, for each jailbreak prompt $p$ and a query set $Q$, the harmfulness scores of all generated responses are computed, and the average of the highest scores is taken as the EMH value, reflecting the most harmful possible responses. In contrast, JSR focuses on the overall success rate of jailbreak prompts, quantifying the probability that a prompt bypasses the LLM’s safety mechanisms. This metric sets a threshold $T$ and calculates the proportion of responses exceeding this threshold to determine the average probability of successful jailbreaks. These two metrics provide complementary perspectives in evaluating LLM jailbreak behaviors: EMH reflects the worst-case potential harm, while JSR measures the average trend of successful jailbreaks. Experimental results indicate a certain degree of positive correlation between the two—prompts that induce detailed harmful responses are often more likely to circumvent safety mechanisms. Additionally, studies find that different jailbreak strategies perform differently in terms of EMH and JSR. Among them, “Virtual AI Simulation” and “Hybrid Strategies” exhibit higher jailbreak success rates across multiple malicious query categories, whereas the “Structured Response” strategy is relatively less effective. Xiao et al. [79] fine-tune a pre-trained DeBERTaV3-large model as a jailbreak detection model and categorize attack success rates into Top-1 ASR and Top-5 ASR. Top-1 ASR measures the success rate of the single best jailbreak template on the target model, while Top-5 ASR calculates the composite success rate of the top five most effective jailbreak templates, that is, if at least one of the five attempts succeeds, it is counted as a success.
V-E2 Jailbreak Evaluation for MLLMs

### v3

@@ IV-A Questions-Answers (Q&A)
First, for question-answering datasets, Gressel et al. [37] introduce a benchmark dataset with 210 prompts sourced from academic literature, Twitter, Medium, and other platforms, emphasizing prompt diversity. It is categorized by “strategies” (8) and “techniques” (33), with each technique having five variants. The dataset includes implicit challenges (165 prompts) and explicit challenges (45 prompts). Röttger et al. [98] propose XSTEST, a test suite of 450 prompts (250 safe and 200 unsafe). The safe prompts are designed to be accepted by well-calibrated models, while the unsafe ones test the model’s ability to reject harmful content.
The StrongREJECT dataset constructed by Souly et al. [100] includes prompts designed to force models into generating harmful content, with an automatic evaluator for measuring jailbreak effectiveness. The evaluator aligns closely with human judgments and reveals that existing evaluation methods overestimate jailbreak success. Souly et al. also find that successful jailbreaks often degrade model performance, highlighting potential impacts on capability. Zhang et al. [95] introduce SafetyBench, a benchmark with 11,435 multiple-choice questions across seven safety categories, supporting both English and Chinese data for bilingual evaluations. Jin et al. [96] create a dataset with 666 jailbreak prompts and 390 harmful questions sourced from high-risk scenarios like hate speech and illegal activities. They curate effective responses and use BERT embeddings to evaluate response effectiveness.
In addition, Gong et al. [70] introduce SafeBench, a safety evaluation benchmark with 500 harmful questions across 10 AI safety topics. The dataset is created in two phases: identifying sensitive topics based on usage policies and generating questions using GPT-4, which are manually filtered for policy violations. They also create SafeBench-Tiny, a smaller subset with 50 questions. Banerjee et al. [102] develop TECHHAZARDQA, a dataset with 7,745 harmful questions spanning seven technical domains. Generated using a fine-tuned Mistral-V2 model, these questions are manually reviewed to ensure they prompt unsafe model behavior. Wang et al. [101] propose a risk taxonomy and compile a dataset of 939 prompts categorized by risk levels, aiming to identify model vulnerabilities. Liu et al. [97] create a training dataset of 1.2 million image-text pairs for academic tasks like visual question answering (VQA), using CLIP-ViT-L-336px for feature extraction and MLP optimization.
@@ IV-B Instructions
For instruction-style datasets, Qiu et al. [99] propose a benchmark to evaluate the safety and robustness of LLMs, introducing a latent jailbreak prompt dataset that embeds malicious instructions within seemingly benign tasks, like translation. They use a hierarchical annotation framework to analyze performance across dimensions such as benign instructions, word substitutions, and instruction placement. SAP (Semi-Automatic Attack Prompts) [88] is a dataset for LLM safety evaluation, containing versions with increasing prompt quantities (SAP5 to SAP200). SAP200 includes 1,600 attack prompts on sensitive topics like fraud, terrorism, and hate speech. The dataset is built through both manual and automated methods, using GPT-3.5-Turbo-0301 to expand it via in-context learning, generating and refining prompts with attack characteristics. This iterative process enhances the dataset’s effectiveness for safety assessment.
Subsequently, to facilitate research on LLM agent misuse, Andriushchenko et al. [42] introduce the AgentHarm benchmark, which includes 110 harmful agent tasks across 11 harm categories, such as fraud, cybercrime, and violence. The dataset, comprising 440 tasks, assesses whether models refuse harmful requests and whether jailbreak agents can still perform multi-step tasks post-attack. JBB-Behaviors [92], part of the JailbreakBench framework, evaluates LLM safety under jailbreak attacks. It includes 100 harmful and 100 benign behaviors across categories like harassment, fraud, and privacy violations. The dataset is generated using various jailbreak methods (PAIR, GCG, JBC) across multiple LLMs, providing a standardized framework for comparing LLM safety and defense capabilities in jailbreak scenarios.
@@ IV-C Harmful Sentences / Images
For datasets in the form of harmful sentences or images, Zou et al. [80] introduce AdvBench, a benchmark designed to evaluate model vulnerability to harmful behaviors. AdvBench consists of 500 harmful strings, covering behaviors such as profanity, violence, threats, misinformation, and cybercrime. These strings, ranging from 3 to 44 tokens, aim to induce models to generate harmful content. Additionally, AdvBench includes 500 instructions targeting similar harmful behaviors, with the goal of finding universal adversarial strings that trigger multiple harmful behaviors. The benchmark has been expanded to 574 harmful strings and 520 instructions through iterative updates. Based on this, Niu et al. [60] categorize harmful behaviors in AdvBench into eight semantic categories (e.g., “bombs or explosives,” “self-harm”) and pair each with semantically relevant images to create AdvBench-M, a multi-modal dataset for evaluating MLLMs’ jailbreak capabilities.
Subsequently, Sun et al. [93] develop SafetyPrompts, a safety evaluation benchmark for Chinese LLMs, which assesses performance across eight safety scenarios and six instruction-based attacks using reward-based prompting. Liu et al. [94] automatically compile a dataset of 5,040 text-image pairs covering 13 harmful scenarios using stable diffusion and layout techniques. Li et al. [33] introduce RTVLM, a dataset containing 750 harmful instructions across five categories (violence, financial crimes, self-harm, privacy violations, animal abuse). The dataset is generated by selecting keywords via GPT-4, creating harmful instructions, and using CLIP ViT-L/14 to select relevant images from Google to enhance attack effectiveness.
@@ V Evaluation Metrics
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=root-line-color, rounded corners, align=center, text centered, minimum width=15em, edge path= [draw=edge-color, line width=1pt] (!parent.east) – +(5pt,0) —- (.child anchor); , s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=1pt, par/.style=rotate=0, child anchor=north, parent anchor=south, anchor=center, minimum height=3em, , where level=1 draw=child-one-line-color, text width=20em, minimum height=3em, font=, , where level=2 draw=child-five-line-color, minimum height=3em, font=, , [ Evaluation Metrics, par, for tree=fill=root-color, [ Human Evaluation $\S$ V-A, for tree=fill=child-one-color, [ xx[36, 31, 35, 30, 89, 73, 75, 79]xx, for tree=fill=child-five-color, ] ] [ Perspective API Evaluation $\S$ V-B, for tree=fill=child-two-color, [ xx[22, 28, 67, 89, 31]xx, for tree=fill=child-five-color, ] ] [ LLM Evaluation $\S$ V-C, for tree=fill=child-three-color, [ xx[26, 25, 22, 31, 34, 33, 35, 28, 89, 29, 81, 88, 37, 73, 90, 40, 79, 83]xx, for tree=fill=child-five-color ] ] [ Keyword Dictionary Evaluation $\S$ V-D, for tree=fill=child-four-color, [ xx[26, 89, 29, 81]xx, for tree=fill=child-five-color ] ] [ Custom Evaluation $\S$ V-E, for tree=fill=child-six-color, [ xx[41, 43, 27, 84, 74, 80, 30, 78, 38, 67, 76, 85, 77, 39]xx, for tree=fill=child-five-color ] ] ] ]
@@ V-A Human Evaluation
This evaluation involves systematic review of model responses by experts or professional annotators, focusing on whether the content exhibits jailbreak behaviors. The assessment includes grading the severity, executability, and potential societal impact of such behaviors. Researchers such as Yu et al. [30] and Shayegani et al. [31] predefine multiple scoring levels and provided systematic training for evaluators. During the evaluation process, each harmful text is independently scored by four experts. If all experts unanimously classify the text as harmful, it is directly categorized as a successful jailbreak case. If three experts deem it harmful, an additional senior reviewer is invited for a final decision to maximize evaluation accuracy and fairness.
While this evaluation mechanism offers significant advantages in evaluation accuracy, it suffers from high costs, poor scalability, and other practical constraints, making it difficult to adapt to large-scale evaluation needs.
@@ V-B Perspective API Evaluation
This evaluation leverages the quantitative analysis capabilities of Perspective API to automatically score text based on toxicity, insultiveness, hatefulness, and other harmful attributes [31, 28]. Specifically, a score above 0.7 indicates the presence of clearly harmful content; scores between 0.3 and 0.7 suggest a high likelihood of harmful information; and scores below 0.3 are generally considered benign.
This evaluation mechanism has shortcomings in cross-language adaptability and contextual understanding, particularly struggling with nuanced language phenomena such as metaphors and sarcasm, which may lead to misclassification. Additionally, for longer texts, Perspective API tends to assign lower harmfulness scores, compromising evaluation accuracy.
@@ V-C LLM Evaluation
This evaluation typically employs high-performance LLMs (e.g., GPT-4, Claude) as the core evaluators, leveraging techniques such as prompt engineering to systematically analyze the harmfulness, policy-violating severity, and jailbreak effectiveness of generated text. For instance, Some researchers [25, 26, 89] instruct GPT-4 to rate the harmfulness of response texts on a scale of 1–10, where 10 indicates explicitly harmful content and is considered a successful jailbreak.
Although this evaluation mechanism demonstrates strong robustness in automated evaluation, in practice, the LLM may refuse to generate ratings for text containing harmful content, compromising the stability and feasibility of the assessment.
@@ V-D Keyword Dictionary Evaluation
This evaluation employs pattern-matching techniques to rapidly detect model outputs based on predefined sensitive word libraries and keyword lists, making it the fastest evaluation approach currently available. Common keyword dictionaries typically include terms such as “I’m sorry”, “Sorry”, “I apologize”, “I’m just”, “can’t”, “I do not”, and “must” [26, 89]. If any of these dictionary terms appear in the model’s response, the jailbreak attempt is considered unsuccessful.
The key challenge lies in constructing a well-designed dictionary that comprehensively covers the model’s refusal patterns to ensure evaluation accuracy and thoroughness.
@@ V-E Custom Evaluation
To address specific research needs, researchers also design tailored evaluation schemes by integrating rule-based matching, contextual analysis, and interactive feedback to construct targeted assessment frameworks. Specifically, we categorize these custom evaluation methods into three groups, depending on whether the jailbreak attacks target LLMs, MLLMs, or Agents.
V-E1 Evaluation for LLMs
@@
Yu et al. [30] introduce two metrics for evaluating jailbreak effectiveness: Expected Maximum Harmfulness (EMH) and Jailbreak Success Rate (JSR). EMH measures the maximum harm of jailbreak inputs under the worst-case scenario by averaging the highest harmfulness scores of generated responses. JSR, on the other hand, calculates the success rate of jailbreak prompts, determining the proportion of responses exceeding a set threshold $T$. These metrics provide complementary views: EMH reflects worst-case harm, while JSR captures the overall success rate. Experimental results show a positive correlation between the two, with prompts inducing harmful responses often bypassing safety mechanisms. Different jailbreak strategies, such as “Virtual AI Simulation” and “Hybrid Strategies,” show higher success rates, while the “Structured Response” strategy is less effective. Xiao et al. [78] fine-tune a DeBERTaV3-large model for jailbreak detection, categorizing attack success rates as Top-1 ASR and Top-5 ASR. Top-1 ASR measures the success of the best jailbreak template, while Top-5 ASR calculates the composite success rate of the top five templates.
V-E2 Evaluation for MLLMs